1/a-b + 1/b-a
I'm confused, why is the answer 0? Could you answer the question in a specific step by step process please, I'd really appreciate it.

Answer by Alan3354(69443)

1/a-b + 1/b-a
-------------
b-a = -(a-b)
---
1/a-b + 1/b-a = 1/a-b - 1/a-b
= 0